OpenStudy (anonymous):

help? Maria and Edwin have $23.00 each to spend at a book fair, where all students receive a 25% discount. They both want to purchase a copy of the same book, which normally sells for $26.50 plus 10% sales tax. To check if she has enough to purchase the book, Maria takes 25% of $26.50 and subtracts that amount from the normal price. She takes 10% of the discounted selling price and adds it back to find the purchase amount. Edwin takes 75% of the normal purchase price and then computes 110% of the reduced price. Is Maria correct? Is Edwin correct? Do they have enough money to purchase the book

OpenStudy (anonymous):

26.50 + 10% -book Maria 23.50 -money 25% - discount what she did: 26.50 - 0.25 (26.50) = 26.50 + (0.25)(26.50) = then add those. That will be the purchase amount of the book that maria calculated. Edwin 23.50 -money 25% - discount what he did: 26.50 x 0.75 x 1.10% = That will be the purchase amount of the book that edwin calculated.

OpenStudy (anonymous):

My dad helped :)

OpenStudy (anonymous):

I hope that answer you queastion

OpenStudy (anonymous):

ur welcome :D

OpenStudy (anonymous):

bye I'll see you around